OpenStudy (anonymous):

a heated piece of metal cools according to the function c(x)=(1/4)^x-8 where x is measured in hours. A device is added that aids in cooling according to the function h(x)=-x-3. What will be the temperature of the metal after five hours?

OpenStudy (anonymous):

@perl @freckles can either of you get this answer?

OpenStudy (anonymous):

i am going to make a guess that you add them up adn get \[c(x)=\left(\frac{1}{4}\right)^{x-8}-x-3\] but it is hard to answer without knowing the original temperature i will make another guess this is a FLVS question

OpenStudy (anonymous):

maybe you just replace \(x\) by \(5\) and compute \[\left(\frac{1}{4}\right)^{5-8}-5-3\]

OpenStudy (anonymous):

you will get some negative number, i think some piece of information might be missing then again with FLVS you never can tell

OpenStudy (anonymous):

this is not from FLVS its from a take home test. I go to public school. Ive found similar questions online it seems like she just changed the numbers.

OpenStudy (anonymous):

ok can you compute that?

OpenStudy (anonymous):

that is, can you compute \[\left(\frac{1}{4}\right)^{5-8}-5-3\]?

OpenStudy (anonymous):

I got 56